# Transportation Market Rate Forecast Using Signature Transform

### v1

@@ Abstract
Currently, Amazon relies on third parties for transportation marketplace rate forecasts, despite the poor quality and lack of interpretability of these forecasts. While transportation marketplace rates are typically very challenging to forecast accurately, we have developed a novel signature-based statistical technique to address these challenges and built a predictive and adaptive model to forecast marketplace rates. This novel technique is based on two key properties of the signature transform. The first is its universal nonlinearity which linearizes the feature space and hence translates the forecasting problem into a linear regression analysis; the second is the signature kernel which allows for comparing computationally efficiently similarities between time series data. Combined, these properties allow for efficient feature generation and more precise identification of seasonality and regime switching in the forecasting process. Preliminary result by the model shows that this new technique leads to far superior forecast accuracy versus commercially available industry models with better interpretability, even during the period of Covid-19 and with the sudden onset of the Ukraine war.
@@ 1 Introduction
Linehaul transportation costs make up a significant portion of overall Amazon transportation costs. To manage these costs, Amazon has developed a variety of tools to manage linehaul capacity mix and procurement. One key input to all of these models is the forecast of transportation marketplace rates, which however are notoriously difficult to forecast – they are driven a number of factors: the ever-changing network of tens of thousands of drivers, shippers of all sizes with a mix of occasional, seasonal, and regular demand, a huge set of brokers, traditional and digital exchanges, and local, regional, national, and international economic factors of all kinds. In addition, the transportation marketplace frequently goes through fundamental shifts – whether because of wars, pandemics, fuel prices, or due to shifting international trade patterns.
Although Amazon has purchased externally-created forecasts for some time, these forecasts are neither explainable nor sufficient/accurate to the specific Amazon needs. To address this challenge, we built a forecasting model based on time series data to predict weekly marketplace rates for the North America market, at both the national and the regional levels. Our approach incorporates an innovative statistical technique capable of efficiently capturing significant fluctuations in transportation marketplace rates.
The key challenge in time series forecasting.
@@
Time series prediction models such as Autoregressive Integrated Moving Average (ARIMA) (Shumway et al. 2017) and Exponential Smoothing (Gardner Jr 2006) assume that the time series are stationary, which is not the case in the marketplace rates. Moreover, ARIMA has limited ability to capture seasonality and long-term trends (Kumar and Vanajakshi 2015), and Exponential Smoothing may be insufficient for abrupt changes or outliers and produce unstable forecasts (Taylor 2004). Furthermore, these methods rely solely on historical data of the time series, which is inadequate in capturing the causal relation between the economic factors and the marketplace rates. Meanwhile, machine learning algorithms such as Long Short-Term Memory Neural Networks (Yu et al. 2019) and Gated Recurrent Units (Chung et al. 2014), though capable of capturing nonlinear relationship and complex patterns in time series data, will require substantially more training data which is not available in our case.
Indeed, one of the main challenges in analyzing time series data is their ever-changing statistical properties, due to factors including changes in business and economic cycles, shifts in policy, or changes in market conditions. In our case, the marketplace itself has recently experienced shifts in regimes and seasonality (Hamilton 1989), in terms of volatility, trends, and cyclical patterns, partly due to the Covid-19 pandemics and the Ukraine-Russian war.
@@
Much of statistical learning theory relies on finding a feature map that embeds the data (for instance, samples of time series) into a high-dimensional feature space. Two requirements for an ideal feature map are universality, meaning that non-linear functions of the data are approximated by linear functionals in the feature space; and characteristicness, meaning that the expected value of the feature map characterizes the law of the random variable. It is shown that with the technique of the signature transform these two properties are in duality and therefore often equivalent (Simon-Gabriel and Schölkopf 2018). This is the main inspiration of our proposed signature-based forecasting technique for our forecast models.
Originally introduced and studied in algebraic topology (Chen 1954, 1957), the signature transform, sometimes referred to as the path signature or simply signature, has been further developed in rough path theory (Lyons et al. 2007, Friz and Victoir 2010), introduced for financial applications (Lyons et al. 2014, Arribas 2018, Lyons et al. 2019, Kalsi et al. 2020) and machine learning (Yang et al. 2015, Xie et al. 2017, Li et al. 2017, Bonnier et al. 2019, Kidger and Lyons 2020), and most recently to time series data analysis (Morrill et al. 2020). Given any continuous or discrete time series, their signature transform produces a vector of real-valued features that extract information such as order and area, and explicitly considers combinations of different channels. The signature of time series uniquely determines the time series, and does so in a computationally efficient way. Most importantly, every continuous function of a time series data may be asymptotically approximated by a linear functional of its signature. In other words, signature transform linearizes the otherwise complicated feature space, and thus is a powerful tool for feature generation and pattern identification in machine learning.
@@
We propose a novel signature-based statistical technique for the time series forecasting problem. This is based on two key properties of the signature transform. The first is the universal nonlinearity of the signature transform which linearizes the features space of the time series data and hence translates the forecasting problem into a linear regression analysis; to avoid issues of overfitting and co-linearity and to improve the forecast accuracy, we adopt the two-step LASSO for the regression analysis (Belloni and Chernozhukov 2013). The second is the signature kernel which allows for computationally efficient comparison of similarities between time series data. Technically, this is to identify different “signature feature maps”, the statistical counterpart of identifying different distributions for a given time series data, albeit in the linearized features space from the signature transform. Combined, this leads to our signature based two-step LASSO approach with adaptive weights via the signature kernel (Chevyrev and Oberhauser 2022). This signature-transform-based technique (Friz and Hairer 2020) for data analysis allows for efficient feature generation and more precise identification of seasonality and regime switching embedded in the data.
Preliminary analysis shows that our forecast model presents superior performance than commercially available forecast models, while providing significantly better interpretability, despite the onset of Covid-19 and the current Ukraine war.
2 Preliminary on Signature Transform
@@ 2.1 Signatures of Continuous Paths
We begin with the definition of signatures of continuous paths.
@@
$=\left(\left(\int_{a<t_{1}<\cdots<t_{k}<b}\mathrm{~{}d}X_{t_{1}}^{i_{1}}\cdots\mathrm{d}X_{t_{k}}^{i_{k}}\right)_{1\leq i_{1},\ldots,i_{k}\leq d}\right)_{k\geq 0},$ (1)
@@
Remark 1
The signature may in fact be defined more generally on paths of bounded variation (Friz and Hairer 2020), but the above definition suffices for our purposes.
Example 1
@@
$\operatorname{Sig}(X)=\left(1,y-x,\frac{1}{2}(y-x)^{\otimes 2},\frac{1}{6}(y-x)^{\otimes 3},\ldots,\frac{1}{k!}(y-x)^{\otimes k},\ldots\right),$ (3)
@@
$\operatorname{Sig}(X)=\left(1,X(b)-X(a),\frac{1}{2}\left(X(b)-X(a)\right)^{2},\frac{1}{6}\left(X(b)-X(a)\right)^{3},\ldots,\frac{1}{k!}\left(X(b)-X(a)\right)^{k},\ldots\right),$ (4)
@@
$\mathbb{E}\left[\operatorname{Sig}(X)\right]=\left(1,\mathbb{E}\left[X(b)-X(a)\right],\frac{1}{2}\mathbb{E}\left[\left(X(b)-X(a)\right)^{2}\right],\ldots,\frac{1}{k!}\mathbb{E}\left[\left(X(b)-X(a)\right)^{k}\right],\ldots\right),$ (5)
whenever it exists, describes precisely the moments of $X(b)-X(a)$. Thus, for a high-dimensional stochastic process $X(t)$, the expected signature, naturally form the generalization of the moments of the process. In other words, for a stochastic process $X(t)$, its expected signature characterizes the law of $X(t)$ up to tree-like equivalence, as proved in Chevyrev and Lyons (2016).
Example 2 shows that the signature for a one-dimensional path, only depends on its total increment. In general, it implies that the signature of a path itself may not carry sufficient information to fully characterize the path. Nevertheless, this problem may be resolved by considering the time-augmented version of the original path.
Definition 2
Given a path $X:[a,b]\rightarrow\mathbb{R}^{d}$, define the corresponding time-augmented path by $\widehat{X}_{t}=\left(t,X_{t}\right)$, which is a path in $\mathbb{R}^{d+1}.$
Theorem 1 (Uniqueness (Hambly and Lyons 2010))
Theorem 2 (Universal nonlinearity (Arribas 2018))
This universal nonlinearity is the key property of the signature transform and important for applications in feature augmentations. See Lyons et al. (2014), Li et al. (2017), Morrill et al. (2020) for examples.
Note that the signature by definition is a infinite dimensional tensor. In practice, one can only compute the truncated signature $\operatorname{Sig}^{N}$ in Equation (2) up to some depth $N$. The next result guarantees that reminder terms in the truncation decay factorially.
Theorem 3 (Factorial decay (Lyons et al. 2007))
Let $X:[a,b]\rightarrow\mathbb{R}^{d}$ be a continuous piecewise smooth path. Then
$\left\|\int_{a<t_{1}<\cdots<t_{k}<b}\mathrm{~{}d}X_{t_{1}}\otimes\cdots\otimes\mathrm{d}X_{t_{k}}\right\|\leq\frac{C(X)^{k}}{k!}$
[…]
The next property about signatures, Theorem 4, is the invariance to time reparameterisations. It implies that the signature encodes the data by its arrival order and independently of its arrival time. This is a desired property in many applications such as hand-writing recognition Yang et al. (2015), Xie et al. (2017). Meanwhile, there is an interesting interplay between Theorem 2 and Theorem 4: In a problem where time parameterisations are irrelevant, it suffices to compute the signature of $X$ by Theorem 4; However, if time parameterisation is important, then according to Theorem 2, applying the signature transform to the time-augmented path $\widehat{X}$ ensures that parameterisation-dependent features are still learned.
Theorem 4 (Invariance to time reparameterisations (Lyons et al. 2007))
2.2 Signatures of Streams of Data
@@
Note that by Theorem 4, the signature of a stream of data is independent of the choice of $u_{i}$ in Equation (6) in a linear interpolation. Meanwhile, by Theorem 2, in order to learn parameterisation-dependent features, one can apply the signature transform to the time-augmented data stream $\widehat{\boldsymbol{x}}=\left(\widehat{\boldsymbol{x}}_{1},\ldots,\widehat{\boldsymbol{x}}_{n}\right)$, where $\widehat{\boldsymbol{x}}_{i}=(\boldsymbol{x}_{i},t_{i})\in\mathbb{R}^{d+1}$, and $t_{i}$ is the time when the data point $\boldsymbol{x}_{i}$ arrives.

### v3

# Transportation Marketplace Rate Forecast Using Signature Transform

@@ Abstract
Freight transportation marketplace rates are typically challenging to forecast accurately. In this work, we have developed a novel statistical technique based on signature transforms and have built a predictive and adaptive model to forecast these marketplace rates. Our technique is based on two key elements of the signature transform: one being its universal nonlinearity property, which linearizes the feature space and hence translates the forecasting problem into linear regression, and the other being the signature kernel, which allows for comparing computationally efficiently similarities between time series data. Combined, it allows for efficient feature generation and precise identification of seasonality and regime switching in the forecasting process.
An algorithm based on our technique has been deployed by Amazon trucking operations, with far superior forecast accuracy and better interpretability versus commercially available industry models, even during the COVID-19 pandemic and the Ukraine conflict. Furthermore, our technique is able to capture the influence of business cycles and the heterogeneity of the marketplace, improving prediction accuracy by more than fivefold, with an estimated annualized saving of \$50 million.
@@ 1 Introduction
Linehaul freight transportation costs make up a significant portion of overall Amazon transportation costs. To manage these costs, Amazon has developed a variety of tools to manage linehaul capacity mix and procurement. One key input to all of these models is a forecast of transportation freight marketplace rates, which however are notoriously difficult to forecast – they are driven by a number of factors: the ever-changing network of tens of thousands of drivers, shippers of all sizes with a mix of occasional, seasonal, and regular demands, a huge set of brokers, traditional and digital exchanges, and local, regional, national, and international economic factors of all kinds. In addition, the transportation marketplace frequently goes through fundamental shifts – either because of wars, pandemics, fuel prices, or due to shifting international trade patterns.
Although Amazon has purchased externally created forecasts for some time, these forecasts are neither explainable nor sufficiently accurate to meet specific Amazon needs. To address this challenge, we have built a forecasting model based on time series data to predict weekly freight marketplace rates for the North America market, at both the national and the regional levels. Our approach incorporates an innovative signature-based statistical technique capable of efficiently capturing significant fluctuations in transportation marketplace rates.
The key challenges in time series forecasting.
@@
Time series prediction models such as Autoregressive Integrated Moving Average (ARIMA) [32] and Exponential Smoothing [14] assume that the time series are stationary, which is not the case for freight marketplace rates. Moreover, ARIMA has limited ability to capture seasonality and long-term trends [21], and Exponential Smoothing may be insufficient for abrupt changes or outliers and produce unstable forecasts [33]. Furthermore, these methods rely solely on historical data from the time series, which is inadequate in capturing the causal relation between the economic factors and the marketplace rates. Meanwhile, machine learning algorithms such as Long Short-Term Memory Neural Networks [37] and Gated Recurrent Units [5], though capable of capturing nonlinear relationship and complex patterns in time series data, will require substantially more training data which is not available in our case.
Indeed, one of the main challenges in analyzing time series data is their ever-changing statistical properties, due to factors including changes in business and economic cycles, shifts in policy, or changes in market conditions. In our case, the market itself has recently experienced shifts in regimes and seasonality [15], in terms of volatility, trends, and cyclical patterns, partly due to the COVID-19 pandemic and the Ukraine conflict.
@@
Much of statistical learning theory relies on finding a feature map that embeds the data (for instance, samples of time series) into a high-dimensional feature space. Two requirements for an ideal feature map are universality, meaning that non-linear functions of the data are approximated by linear functionals in the feature space; and characteristicness, meaning that the expected value of the feature map characterizes the law of the random variable. It is shown that with the technique of the signature transform these two properties are in duality and therefore often equivalent [31]. This is the primary inspiration for our proposed signature-based forecasting technique for our forecast models.
Originally introduced and studied in algebraic topology [6, 7], the signature transform, sometimes referred to as the path signature or simply signature, has been further developed in rough path theory [22, 12], introduced for financial applications [24, 1, 25, 19] and machine learning [36, 35, 26, 3, 18], and most recently to time series data analysis [27, 10, 11]. Given any continuous or discrete time series, their signature transform produces a vector of real-valued features that extract information such as order and area, and explicitly considers combinations of different channels. The signature of time series uniquely determines the time series, and does so in a computationally efficient way. Most importantly, every continuous function of a time series data may be asymptotically approximated by a linear functional of its signature. In other words, signature transform linearizes the otherwise complicated feature space, and thus is a powerful tool for feature generation and pattern identification in machine learning.
@@
We propose a novel signature-based statistical technique for the time series forecasting problem. This is based on two key elements of the signature transform. The first is the universal nonlinearity property of the signature transform, which linearizes the feature space of the time series data and hence translates the forecasting problem into a linear regression. The second is the signature kernel which allows for computationally efficient comparison of similarities between time series data. Technically, this is to identify different “signature feature maps”, the statistical counterpart of identifying different distributions for a given time series data, albeit in the linearized feature space from the signature transform.
Our approach starts by collecting data including a hundred of market supply and demand factors, and runs a correlation test between the marketplace rate and the factors to remove non-significant factors and to identify factors that may be colinear. We then exploit the universal nonlinearity property of signature transform to construct signature features as suitable candidates for the “internal” features. To avoid issues of overfitting and co-linearity between the signature feature and the external factors, and to improve the forecast accuracy, we adopt the two-step LASSO for the regression analysis [2]. Finally, this two-step LASSO is enhanced with adaptive weight using a signature kernel, which enables capturing changes in regimes or seasonality. Combined, this leads to our signature-based adaptive two-step LASSO approach. This novel signature-transform-based technique for data analysis allows for efficient feature generation and more precise identification of seasonality and regime switching embedded in the data.
Implementation and real-time performance.
This signature-based adaptive two-step LASSO algorithm has been implemented for the trucking operations in Amazon since November 2022. Performance analysis shows that our forecast model presents superior performance than commercially available forecast models, while providing significantly better interpretability. Despite the onset of COVID-19 and the Ukraine conflict, it captures the influence of business cycles and heterogeneity of the marketplace, improves prediction accuracy by more than fivefold, and has an estimated annualized saving of approximately \$$\$50$ million.
2 Technical Background
@@ 2.1 Signatures of Continuous Paths
We begin with the definition of signatures of continuous piecewise smooth paths.
@@
$=\left(\left(\int_{a<t_{1}<\cdots<t_{k}<b}\mathrm{~{}d}X_{t_{1}}^{i_{1}}\cdots\mathrm{d}X_{t_{k}}^{i_{k}}\right)_{1\leq i_{1},\ldots,i_{k}\leq d}\right)_{k\geq 0},$
@@
Remark 1.
The signature can be defined more generally on paths of bounded variation [11], but the above definition suffices for our purposes.
Example 1.
@@
$\operatorname{Sig}(X)=\left(1,y-x,\frac{1}{2}(y-x)^{\otimes 2},\frac{1}{6}(y-x)^{\otimes 3},\ldots,\frac{1}{k!}(y-x)^{\otimes k},\ldots\right),$
@@
$\operatorname{Sig}(X)=\Big{(}$ $1,X(b)-X(a),\frac{1}{2}\left(X(b)-X(a)\right)^{2},\frac{1}{6}\left(X(b)-X(a)\right)^{3},$
$\quad\ldots,\frac{1}{k!}\left(X(b)-X(a)\right)^{k},\ldots\Big{)},$
@@
$\mathbb{E}\left[\operatorname{Sig}(X)\right]=\Big{(}$ $1,\mathbb{E}\left[X(b)-X(a)\right],\frac{1}{2}\mathbb{E}\left[\left(X(b)-X(a)\right)^{2}\right],$
$\quad\ldots,\frac{1}{k!}\mathbb{E}\left[\left(X(b)-X(a)\right)^{k}\right],\ldots\Big{)},$
whenever it exists, describes precisely the moments of $X(b)-X(a)$. Thus, for a high-dimensional stochastic process $X(t)$, the expected signature naturally forms the generalization of the moments of the process. In other words, for a stochastic process $X(t)$, its expected signature characterizes the law of $X(t)$ up to tree-like equivalence, as proved in [8].
Example 2 shows that the signature for a one-dimensional path only depends on its total increment. In general, it implies that the signature of a path itself may not carry sufficient information to fully characterize the path. Nevertheless, this problem may be resolved by considering the time-augmented version of the original path (see Definition 4, Theorem 1 and 2 below).
2.2 Signature of Discrete Data
@@
Definition 4.
Given a path $X:[a,b]\rightarrow\mathbb{R}^{d}$, define the corresponding time-augmented path by $\widehat{X}_{t}=\left(t,X_{t}\right)$, a path in $\mathbb{R}^{d+1}.$
2.3 Key Properties of Signature
Theorem 1 (Uniqueness [16]).
Theorem 2 (Universal nonlinearity [1]).
This universal nonlinearity is the key property of the signature transform and is important for our model, and in general for applications in feature augmentations. See [24], [26], [27] for examples.
Note that the signature by definition is an infinite-dimensional tensor. In practice, one can only compute the truncated signature $\operatorname{Sig}^{N}$ in (1) up to some depth $N$. The next result guarantees that reminder terms in the truncation decay factorially.
Theorem 3 (Factorial decay [22]).
[…]
$\left\|\int_{a<t_{1}<\cdots<t_{k}<b}\mathrm{~{}d}X_{t_{1}}\otimes\cdots\otimes\mathrm{d}X_{t_{k}}\right\|\leq\frac{C(X)^{k}}{k!},$
where $C(X)$ is a constant depending on $X$.
The next property about signatures, Theorem 4, is the invariance to time reparameterizations. It implies that the signature encodes the data by its arrival order and independently of its arrival time. This is a desired property in many applications such as hand-writing recognition [36], [35]. Meanwhile, there is an interesting interplay between Theorem 2 and Theorem 4: in a problem where time parameterizations are irrelevant, it suffices to compute the signature of $X$ by Theorem 4; However, if time parameterization is important, then according to Theorem 2, applying the signature transform to the time-augmented path $\widehat{X}$ ensures that parameterization-dependent features are still learned.
Theorem 4 (Invariance to time reparameterizations [22]).
Note that by Theorem 4, the signature of a stream of data is independent of the choice of $u_{i}$ in a linear interpolation in (2). Meanwhile, by Theorem 2, in order to learn parameterization-dependent features, one can apply the signature transform to the time-augmented data stream $\widehat{\boldsymbol{x}}=\left(\widehat{\boldsymbol{x}}_{1},\ldots,\widehat{\boldsymbol{x}}_{n}\right)$, where $\widehat{\boldsymbol{x}}_{i}=(t_{i},\boldsymbol{x}_{i})\in\mathbb{R}^{d+1}$, and $t_{i}$ is the time when the data point $\boldsymbol{x}_{i}$ arrives.